Dawson Toys, Ltd., produces a toy called the Maze. The company has recently established a standard cost system to help control costs and has established the following standards for the Maze toy:

Direct materials: 7 microns per toy at $0.32 per micron
Direct labor: 1.4 hours per toy at $6.70 per hour
During July, the company produced 5,400 Maze toys. Production data for the month on the toy follow:
Direct materials: 72,000 microns were purchased at a cost of $0.31 per micron. 24,750 of these microns were still in inventory at the end of the month.
Direct labor: 8,060 direct labor-hours were worked at a cost of $59,644.
Required:
1.

Compute the following variances for July: (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Do not round intermediate calculations. Round final answer to the nearest whole dollar.)

a. The materials price and quantity variances.

         

b. The labor rate and efficiency variances.
Answer #1

Material price variance = (standard price - actual price)*actual quantity

=(0.32- 0.31)*72,000

=$720 F

Matera quantity variance = (standard quantity - actual quantity used)*standard price

=(5,400*7-72,000)*0.32

=$3,024 U

Labor rate variance =(standard rate- actual rate)*actual hours

=(6.70- actual rate)*8,060

=6.70*8060 - 59,644

=$5,642 U

Labor efficiency variance =(standard hours - actual hours)*standard rate

=(5,400*1.4 - 8,060)*6.7

=$3,350 U